# Automated generation of photonic circuits for Bell tests with homodyne measurements

### v1

@@ 1 Introduction
The nonlocality certified by a violation of the Clauser-Horne-Shimony-Holt (CHSH) inequality [15], a bipartite Bell inequality, stands at the core of numerous DI protocols. Loophole-free violations of the CHSH inequality have been accomplished using a variety of platforms, from NV-centers [16] and photonic setup [17, 18, 19, 20, 21] to neutral atoms [22] and superconducting circuits [23]. Efforts aiming to generate high CHSH violations at high rates have culminated with the first distribution of a device-independent key using trapped ions [24]. Results have also been obtained to extend DIQKD over hundreds of meters with single atoms [25]. On-going efforts [26] aims at implementing device-independent quantum information processing with a purely photonic platform – a platform where optical modes are entangled, manipulated and detected – which is plausibly closer to what is expected for a commercial device.
With their high efficiency, low noise, high bandwidth and capability to operate at telecom wavelengths at room temperature, homodyne detectors are a natural candidate for implementing practical Bell tests. Coupled with a fully photonic circuit for state generation and manipulation, Bell tests may be implemented using integrated on-chip devices. Encouragingly, it has been shown that high CHSH violations can be achieved with homodyne measurements [27, 28, 29, 30]. However, these proposals require tailored states whose experimental realizations is either unknown or, in the best of cases, very challenging. Meanwhile, realistic optical setups with homodyne detectors achieving a CHSH score of $\mathcal{B}\approx 2.048$ have been proposed [31, 32]. This score is, however, below the threshold required by some DI protocols [33, 34]. Therefore, designing realistic optical setups with homodyne detection leading to higher CHSH scores remains an open problem.
@@
In this work, we tackle circuit discovery with an automated approach that combines machine learning with an efficient simulation of optical circuits based on Gaussian representation of quantum states. Specifically, we use deep reinforcement learning for the design of practical photonic CHSH tests, characterized by standard optical components and homodyne measurements. Deep reinforcement learning has proven to be highly effective in discovering relevant elements in a large exploration space [37, 38, 39, 40, 41] and has already been applied to many quantum physics tasks [42, 43, 44, 45], including optical circuits design [46, 47, 48, 49].
@@ 2.1 The CHSH Inequality
=|$\mathcal{B}=\absolutevalue{\langle{A_{0}B_{0}}\rangle+\langle{A_{0}B_{1}}\rangle+\langle{A_{1}B_{0}}\rangle-\langle{A_{1}B_{1}}\rangle}$ | (2)
@@ 2.2 Photonic Circuits
As we focus on practical photonic implementations of Bell tests, we first revisit the relevant concepts of quantum optics. After a brief introduction to photonics, we present Gaussian states and operations that can be accessed using standard optical components. We then review homodyne measurements and the binning of their outcomes, which are used to compute the CHSH score. Finally, we highlight the heralding of quantum states based on detection events with threshold detectors. Further details are reported in Appendix A.
@@
The quadrature field operators of a bosonic mode are expressed in terms of ladder operators using $\hat{x}=(\hat{a}^{\dagger}+\hat{a})/2$ and $\hat{p}=i(\hat{a}^{\dagger}-\hat{a})/2$, such that $[\hat{x},\hat{p}]=i/2$ holds. Given $N$ bosonic modes, we arrange these operators in a vector $\mathbf{\hat{q}}=(\hat{x}_{1},\hat{p}_{1},\ldots,\hat{x}_{n},\hat{p}_{n})$.
@@
where $M$ is a symplectic matrix. In this manuscript we consider four types of gates: two passive operations, beam splitters, $\hat{B}(\theta)$ and phase shifters $\hat{R}(\theta)$, as well as two active ones, namely single-mode and two-mode squeezers, labeled $\hat{S}_{1}(r)$ and $\hat{S}_{2}(r)$ respectively. Note that these operations are all characterized by $\mathbf{d}=\vec{0}$, therefore leaving the displacement vector invariant if applied to a state with ${\bm{\mu}}=\vec{0}$. The exact expressions of these transformations are reported in Table 1 of Section A.2.
@@
for the correlators defined in Eq. 1. In Section A.4, we detail how we numerically compute the above expression.
@@
To obtain states with negative Wigner functions, we herald Gaussian states using threshold detectors. In particular, a heralded two-mode state $\rho_{12}$ can be obtained from an $N$-mode Gaussian state by measuring the presence of one (or more) photons in the other $N-2$ modes [54]. The resulting state can be written as a linear combination of Gaussian states
@@
with coefficients $\omega_{k}\in\mathbb{R}$ satisfying  [^1] Note that the coefficients $\omega_{k}$ need to satisfy additional constraints in order to ensure the positivity of the state $\hat{\rho}_{12}$. $\sum_{k}\omega_{k}=1$, and where $\hat{\rho}^{k}_{12}$ are normalized two-mode Gaussian states. Since the Wigner function is linear in the density operator, the Wigner function of such heralded state may become negative in part of the phase space when some of the coefficients $\omega_{k}$ are negative. Further details on the expression of heralded states and their creation with threshold detectors are given in Section A.3.
@@ 3.1 Photonic Bell tests as a learning task
In the following, we formulate the tasks of finding photonic Bell tests in the agent-environment framework of reinforcement learning. More specifically, we present the environment implementations and the set of actions the agent can interact with, the state and reward received as feedback, as well as an episode routine from initial conditions to termination.
The environment is chosen to be an $N$-mode optical circuit, where the first mode is sent to Alice while Bob receives the second one. As an initial condition, each mode is set to the vacuum state. The agent acts on the environment by appending photonic gates to the optical circuit. Each action is identified by the selected gate, either a beam splitter $\hat{B}(\theta)$, phase shifter $\hat{R}(\theta)$, single-mode squeezer $\hat{S}_{1}(r)$ or two-mode squeezer $\hat{S}_{2}(r)$, and the mode or modes on which the gate is applied. At this stage, the parameters of each photonic gate are fixed to a predefined value in order to limit the size of the action space during training. After each action on the environment, the agent receives in return a state and a reward. Finally, once the circuit depth (i.e. number of gates) reaches a maximum predefined value $n_{\rm circuit}$, we terminate the episode and reset the photonic circuit to empty modes.
The state returned by the environment is given by the photonic state shared between Alice and Bob. This state is computed by applying the current gates in the circuit and by heralding the state of the first two modes on a photon detection in each of the last $N-2$ modes. This computation is efficiently carried out numerically by the code which is made available in Ref. [55]. As the heralded state can be seen as a linear combination of $2^{N-2}$ Gaussian states, as in Eq. 10, we represent it as a vector of $2^{N-2}(10+1)$ real elements. More specifically, the vector is a concatenation over all $k$ of the $10$ elements of the covariance matrix of each Gaussian state, c.f. Section A.1, $\hat{\rho}_{12}^{k}$, followed by the coefficients $\omega_{k}$ of the linear combination (remember that the Gaussian parameter $\mu$ is always zero here so we do not need to store it). Note that if the photon detection probability in one of the heralded modes is below a certain threshold, set to $10^{-10}$, the heralding is considered failed to avoid numerical errors (note that this still allows for a global heralding probability smaller than $10^{-10}$ in the presence of several heralding modes). In this case, the state passed to the agent is a trivial two-mode vacuum state.
@@
This expression of the reward has been crafted empirically to favour high CHSH score, even when small improvement are observed.
@@ 3.2 Learning to design Bell tests
A PPO agent is entrusted with learning to design practical photonic Bell tests. This makes use of a two-headed four-layer neural network, for the policy and value network. The input layer encodes the parameters of the heralded state while the neurons of the output layer represent different actions for the policy head, and the single neuron of the value-head returns the scalar $V(s)$. The number of neurons in the two hidden layers is chosen depending on the size of the state parametrization and the size of the action space. For example, we choose $150$ and $90$ neurons in the first and second hidden layer, respectively, when testing circuits with $N=6$ modes where we restrict the agent’s actions to only passive gates. We instead choose 45 and 30 neurons for $N=4$ modes and arbitrary gates. We use the PPO agent implementation of the Julia Reinforcement Learning package [56], and rely on the default hyperparameters except for the update frequency and the trajectory capacity, which are set to 64 and 256, respectively.
@@ 3.3 Exploration and learning strategies
Concretely, we consider four supplementary strategies in which the circuit is initialized with some fixed active gates that generate photons (single- or two-mode squeezers) and the agent is only allowed to choose gates that are passive (phase shifter and beam splitter) for the rest of the episode. The specific initialization of each strategy is described in Appendix B. These new strategies avoid the unnecessary action choices present in the first strategy because the circuit is initialized in a state containing photons in each mode. Additionally, the risk of too large squeezing parameters is reduced by the presence of at most one squeezer per mode. This constraint also favors experimentally realistic setups as chaining squeezing operations appears challenging. It should be noted that despite limiting the agent to passive gates only, these strategies are not restrictive. Indeed, the Euler decomposition (c.f. Appendix A) guarantees that an arbitrary circuit initialized in the vacuum can always be reproduced by a circuit initialized with squeezers and completed by only passive gates. Finally, by reducing the size of the action space, such strategies can favour the learning process of the agent.
In order to identify the most suitable of these exploration strategies, we test them on random circuits. Namely, we construct circuits by appending successively uniformly chosen random gates from the chosen set until the circuit depth $n_{\mathrm{circuit}}$ is reached. With no reliance on neural networks, such a random search is light on computational resource and, therefore, allows for the automated exploration of many more circuits in a given time. In particular, a random exploration may provide a circuit yielding a suitable CHSH score. This exploration is however unstructured and less scalable. In our case, we use the result of this investigation to identify the most suitable exploration strategies for reinforcement learning (see Appendix B for more details).
@@
Finally, for each considered circuit building strategy, we consider two heralding mechanisms for the last $N-2$ modes. In the first scheme, we select states based on detection events from threshold detectors applied to each of the last $N-2$ modes. A threshold detector differentiates between vacuum and the presence of one or more photons. In the second heralding scheme, we aim to better approximate the single-photon subtraction operation, which is known to introduce non-Gaussianity that is essential for generating nonlocality [54]. This is achieved using an unbalanced beam splitter with high transmittivity, along with a threshold detector at each output port. We select events where a click occurs in the reflected port and a vacuum is detected in the transmitted port. For further details on both heralding schemes, we refer the reader to Appendix B.
@@ 4 Results
In this section we discuss the optical circuits for CHSH violation obtained with our automated approach. We then focus on a specific circuit, depicted in Fig. 1, which has the merit of being relatively simple while providing a significant CHSH score. We analyse the resilience of this circuit to photon loss with distance between the parties, as well as its compatibility with inefficient detectors.
@@
However, we also observe that in a given allocated time, the agent could not consistently outperform the random search. In presence of $N=6$ modes, for instance, the best circuit identified by the PPO approach reached $\mathcal{B}=2.072$, whereas random search achieved $\mathcal{B}=2.076$. It is worth noting that the corresponding circuit is significantly more convoluted, though, involving 42 gates for the random search compared to just 8 gates for the solution provided by the PPO agent. This non-optimal CHSH score is partially due to the fact that our agent is trained with modest computing power and does not rely on a large neural network. Therefore, the agent cannot learn complex circuit patterns required to reach higher CHSH scores. Moreover, learning a suboptimal policy has the additional negative effect of reducing the exploration in later episodes, which further reduces the chances to find highly non-trivial circuits leading to higher CHSH scores. A summary of the circuits found by the PPO agent and by random search for the different strategies is provided in Appendix B.
We now focus on the circuit depicted in Fig. 1. This circuit appears as a good candidate for implementing a photonic Bell test with homodyne measurements. It has been found by the PPO agent interacting with a circuit of $N=4$ modes initialized with two two-mode squeezers and is composed of only two additional beam splitters. The third and fourth modes are used to herald the state using threshold detectors. This photonic circuit yields a CHSH score of $\mathcal{B}=2.068$ with optimal squeezing parameters corresponding to $3.9$ dB and $8\cdot 10^{-3}$ dB of squeezing, respectively. The heralding probability of the state is $3\cdot 10^{-6}$. Considering squeezers generating pulses at the MHz rate, the proposed circuit can prepare about $300$ states per second. Therefore, the statistics for a Bell test can be gathered in a realistic amount of time, e.g. $10^{6}$ shots in an hour.
Figure 2: CHSH score with respect to the distance between Alice, where state preparation occurs, and Bob. We consider an optical fiber with a loss of $0.2$ dB/km, and numerically optimized the circuits’ parameters for each step of $0.1$ km. As higher squeezing is beneficial for greater distance, we ensure squeezing of at most $10dB$ to respect practical experimental range.
@@
To study the effect of photon loss, we assume that the optical circuit preparing and heralding the quantum state is located in Alice’s laboratory. Bob receives his part of the state, i.e the second mode, through an optical fiber link characterized by $0.2$ dB/km of attenuation. We assume that the optical signals are prepared at telecom wavelength, and, hence, that no conversion are necessary. In Fig. 2, we plot the CHSH score as a function of the distance between Alice and Bob. For every value of distance, we re-optimize the gates parameters to maximize the CHSH score. We observe that a CHSH violation is still possible for distances up to $8$ km. As for the heralding probability of the state, this is not affected by photon loss occurring in Alice’s or Bob’s modes.
We then analyse the effect of imperfect detectors on circuit performance. Specifically, we assume that the threshold detectors used in the heralded modes have efficiency $\eta<1$. In Fig. 3 we plot the heralding probability as a function of $\eta$, for fixed gate parameters optimal in the $\eta=1$ case. We observe that the heralding probability decreases by at most one order of magnitude for efficiencies $\eta\geq 0.25$. Moreover, we observe that the efficiency of the detectors has barely any impact on the CHSH score. Specifically, from a value $\mathcal{B}=2.068$ at $\eta=1.0$, the CHSH score approaches $\mathcal{B}=2.067$ as $\eta\rightarrow 0$. This can be explained by the fact that the average number of photons reaching the threshold detectors is well below $1$. Indeed, if an inefficient threshold detector measures one mode out of a multi-photon entangled state, it detects each $n$-photon component of the state with a different probability, $1-(1-\eta)^{n}$, thereby modifying the photon number distribution of the unmeasured modes accordingly. This causes the conditional state of the unmeasured modes to depend, in principle, on the detector’s efficiency. However, if the multi-photon components in the measured mode are suppressed, the dependence of the conditional state on the detector efficiency becomes negligible, and the only real consequence is a reduction in the heralding probability.

### v2

@@ 1 Introduction
The nonlocality certified by a violation of the Clauser-Horne-Shimony-Holt (CHSH) inequality [15], a bipartite Bell inequality, stands at the core of numerous DI protocols. Loophole-free violations of the CHSH inequality have been accomplished using a variety of platforms, from NV-centers [16] and photonic setup [17, 18, 19, 20, 21] to neutral atoms [22] and superconducting circuits [23]. Efforts aiming to generate high CHSH violations at high rates have culminated with the first distribution of a device-independent key using trapped ions [24]. Results have also been obtained to extend DIQKD over hundreds of meters with single atoms [25]. On-going efforts [26] aim at implementing device-independent quantum information processing with a purely photonic platform – a platform where optical modes are entangled, manipulated and detected – which is plausibly closer to what is expected for a commercial device.
With their high efficiency, low noise, high bandwidth and capability to operate at telecom wavelengths at room temperature, homodyne detectors are a natural candidate for implementing practical Bell tests. Coupled with a fully photonic circuit for state generation and manipulation, Bell tests may be implemented using integrated on-chip devices. Encouragingly, it has been shown that high or even maximal CHSH violations can be achieved with homodyne measurements [27, 28, 29, 30]. However, these proposals require tailored states whose experimental realizations is either unknown or, in the best of cases, very challenging. Specifically, the proposal of the physical process required to implement the state yielding a maximal CHSH score $\mathcal{B}$ proposed in [29] is very far fetched. Likewise, the state producing a violation of $\mathcal{B}=2.14$ in [30] has a challenging implementation requiring six bosonic modes, each with high squeezing value, and four single-photon detectors. Meanwhile, the best known CHSH score based on a realistic optical setup, which uses standard optics equipment and homodyne detectors, achieves a CHSH score of $\mathcal{B}\approx 2.048$ [31, 32]. This score is, however, below the threshold required by some DI protocols. A notable example is self-testing the singlet state, which is infeasible for CHSH scores less than $2.051$ [33], while the most advanced protocol requires a score $\gtrapprox 2.106$ [34]. Therefore, designing realistic optical setups with homodyne detection leading to higher CHSH scores remains an open problem.
@@
In this work, we tackle circuit discovery with an automated approach that combines a search algorithm with an efficient simulation of optical circuits based on the Gaussian representation of quantum states. Specifically, we consider the design of practical photonic CHSH tests, characterized by standard optical components and homodyne measurements. Among the search strategies we employ, one is based on deep reinforcement learning, which has proven highly effective in identifying relevant elements in large and complex exploration spaces [37, 38, 39, 40, 41] and has already been successfully applied to a variety of quantum physics tasks [42, 43, 44, 45], including the design of optical circuits [46, 47, 48, 49]. We also contrast performances of reinforcement learning with that of a simple random search over the same circuit space.
@@ 2.1 The CHSH Inequality
$|\mathcal{B}=\absolutevalue{\langle{A_{0}B_{0}}\rangle+\langle{A_{0}B_{1}}\rangle+\langle{A_{1}B_{0}}\rangle-\langle{A_{1}B_{1}}\rangle}$ (2)
@@ 2.2 Photonic Circuits
As we focus on practical photonic implementations of Bell tests, we first revisit the relevant concepts of quantum optics. After a brief introduction to photonics, we present Gaussian states and operations that can be accessed using standard optical components. We then review homodyne measurements and the binning of their outcomes, which are used to compute the CHSH score. Finally, we highlight the heralding of quantum states based on detection events with threshold detectors. Further details are reported in Appendix˜A.
@@
The quadrature field operators of a bosonic mode are expressed in terms of ladder operators using $\hat{x}=(\hat{a}^{\dagger}+\hat{a})/2$ and $\hat{p}=i(\hat{a}^{\dagger}-\hat{a})/2$, such that $\hat{x},\hat{p}]=i/2$ holds. Given $N$ bosonic modes, we arrange these operators in a vector $\mathbf{\hat{q}}=(\hat{x}_{1},\hat{p}_{1},\ldots,\hat{x}_{n},\hat{p}_{n})$.
@@
where $M$ is a symplectic matrix. In this manuscript we consider four types of gates: two passive operations, beam splitters, $\hat{B}(\theta)$ and phase shifters $\hat{R}(\theta)$, as well as two active ones, namely single-mode and two-mode squeezers, labeled $\hat{S}_{1}(r)$ and $\hat{S}_{2}(r)$ respectively. Note that these operations are all characterized by $\mathbf{d}=\vec{0}$, therefore leaving the displacement vector invariant if applied to a state with ${\bm{\mu}}=\vec{0}$. The exact expressions of these transformations are reported in Table˜1 of Section˜A.2.
@@
for the correlators defined in Eq.˜1. In Section˜A.4, we detail how we numerically compute the above expression.
@@
To obtain states with negative Wigner functions, we herald Gaussian states using threshold detectors. In particular, a heralded two-mode state $\rho_{12}$ can be obtained from an $N$-mode Gaussian state by measuring the presence of one (or more) photons in the other N2N-2 modes [54]. The resulting state can be written as a linear combination of Gaussian states
@@
with coefficients $\omega_{k}\in\mathbb{R}$ satisfying [^1] Note that the coefficients $\omega_{k}$ need to satisfy additional constraints in order to ensure the positivity of the state $\hat{\rho}_{12}$. $\sum_{k}\omega_{k}=1$, and where $\hat{\rho}^{k}_{12}$ are normalized two-mode Gaussian states. Since the Wigner function is linear in the density operator, the Wigner function of such heralded state may become negative in part of the phase space when some of the coefficients $\omega_{k}$ are negative. Further details on the expression of heralded states and their creation with threshold detectors are given in Section˜A.3.
@@ 3 Automated generation of photonic Bell tests
We provide the full codebase implementing the environment, agent, and circuit-exploration strategies described in this section in an online repository [55].
@@ 3.1 Photonic Bell tests as a learning task
In the following, we formulate the tasks of finding photonic Bell tests in the agent-environment framework. More specifically, we present the environment implementations and the set of actions the agent can interact with, the state and reward received as feedback, as well as an episode routine from initial conditions to termination. Note that this framework is not restricted to reinforcement learning agents and can be used by other search algorithms.
The environment is chosen to be an $N$-mode optical circuit, where the first mode is sent to Alice and the second to Bob, while the remaining N2N-2 modes are heralded during state preparation. As an initial condition, each mode is set to the vacuum state. The agent acts on the environment by appending photonic gates to the optical circuit. Each action is identified by the selected gate, either a beam splitter $\hat{B}(\theta)$, phase shifter $\hat{R}(\theta)$, single-mode squeezer $\hat{S}_{1}(r)$ or two-mode squeezer $\hat{S}_{2}(r)$, and the mode or modes on which the gate is applied. At this stage, the parameters of each photonic gate are fixed to a predefined value in order to limit the size of the action space during training. After each action on the environment, the agent receives in return a state and a reward. Finally, once the circuit depth (i.e. number of gates) reaches a maximum predefined value $n_{\rm circuit}$, we terminate the episode and reset the photonic circuit to empty modes.
The state returned by the environment is given by the photonic state shared between Alice and Bob. This state is computed by applying the current gates in the circuit and by heralding the state of the first two modes on a photon detection in each of the last N2N-2 modes. This computation is efficiently carried out numerically by the code which is made available in Ref. [56]. As the heralded state can be seen as a linear combination of $2^{N-2}$ Gaussian states, as in Eq.˜10, we represent it as a vector of $2^{N-2}(10+1)$ real elements. More specifically, the vector is a concatenation over all $k$ of the $10$ elements of the covariance matrix of each Gaussian state, c.f. Section˜A.1, $\hat{\rho}_{12}^{k}$, followed by the coefficients $\omega_{k}$ of the linear combination (remember that the Gaussian parameter $\mu$ is always zero here so we do not need to store it). Note that if the photon detection probability in one of the heralded modes is below a certain threshold, set to $10^{-10}$, the heralding is considered failed to avoid numerical errors (note that this still allows for a global heralding probability smaller than $10^{-10}$ in the presence of several heralding modes). In this case, the state passed to the agent is a trivial two-mode vacuum state.
@@
This reward function was designed empirically to bias the agent toward circuits that violates the CHSH inequality. In particular, non-violating circuits, i.e. $\mathcal{B}<2$, receive only a weak, linear reward, whereas circuits achieving a violation are assigned a reward that grows exponentially with the CHSH score. This exponential scaling encourages the agent not only to discover violating circuits, but also to favor improvements in the score, even when such improvements are marginal.
@@ 3.2 Learning to design Bell tests
A PPO agent is entrusted with learning to design practical photonic Bell tests. This makes use of a two-headed four-layer neural network, for the policy and value network. The input layer encodes the parameters of the heralded state while the neurons of the output layer represent different actions for the policy head, and the single neuron of the value-head returns the scalar $V(s)$. The number of neurons in the two hidden layers is chosen depending on the size of the state parametrization and the size of the action space. For example, we choose $150$ and $90$ neurons in the first and second hidden layer, respectively, when testing circuits with $N=6$ modes where we restrict the agent’s actions to only passive gates. We instead choose 45 and 30 neurons for $N=4$ modes and arbitrary gates. We use the PPO agent implementation of the Julia Reinforcement Learning package [57], and rely on the default hyperparameters, found in Table˜3, except for the update frequency and the trajectory capacity, which are set to 64 and 256, respectively.
@@ 3.3 Exploration and learning strategies
Concretely, we consider four supplementary strategies in which the circuit is initialized with some fixed active gates that generate photons (single- or two-mode squeezers) and the agent is only allowed to choose gates that are passive (phase shifter and beam splitter) for the rest of the episode. The specific initialization of each strategy is described in Appendix˜B. These new strategies avoid the unnecessary action choices present in the first strategy because the circuit is initialized in a state containing photons in each mode. Additionally, the risk of too large squeezing parameters is reduced by the presence of at most one squeezer per mode. This constraint also favors experimentally realistic setups as chaining squeezing operations appears challenging. It should be noted that despite limiting the agent to passive gates only, these strategies are not restrictive. Indeed, the Euler decomposition (c.f. Appendix˜A) guarantees that an arbitrary circuit initialized in the vacuum can always be reproduced by a circuit initialized with squeezers and completed by only passive gates. Finally, by reducing the size of the action space, such strategies can favour the learning process of the agent.
In order to identify the most suitable of these exploration strategies, we test them on random circuits. Namely, we construct circuits by appending successively uniformly chosen random gates from the chosen set until the circuit depth $n_{\mathrm{circuit}}$ is reached. With no reliance on neural networks, such a random search is light on computational resource and, therefore, allows for the automated exploration of many more circuits in a given time. In particular, a random exploration may provide a circuit yielding a suitable CHSH score. This exploration is however unstructured and may become less effective as the size of the search space grows. In our case, we use the result of this investigation to identify the most suitable exploration strategies for reinforcement learning (see Appendix B for more details).
@@
Finally, for each considered circuit building strategy is explored using each of two heralding mechanisms for the last N2N-2 modes. In the first scheme, we select states based on detection events from threshold detectors applied to each of the last N2N-2 modes. A threshold detector differentiates between vacuum and the presence of one or more photons. In the second heralding scheme, we aim to better approximate the single-photon subtraction operation, which is known to introduce non-Gaussianity that is essential for generating nonlocality [54]. This is achieved using an unbalanced beam splitter with high transmittivity, along with a threshold detector at each output port. We select events where a click occurs in the reflected port and a vacuum is detected in the transmitted port. For further details on both heralding schemes, we refer the reader to Appendix˜B.
@@ 4 Results
In this section we discuss the optical circuits for CHSH violation obtained with our automated approach. We then focus on a specific circuit, depicted in Fig.˜1, which has the merit of being relatively simple while providing a significant CHSH score. We analyse the resilience of this circuit to photon loss with distance between the parties, as well as its compatibility with inefficient detectors. To ensure reproducibility, the Julia notebook used to produce the analysis and plots of this section can be found online [55].
@@
However, we also observe that in a given allocated time, the agent could not consistently outperform the random search. In presence of $N=6$ modes, for instance, the best circuit identified by the PPO approach reached $\mathcal{B}=2.072$, whereas random search achieved $\mathcal{B}=2.076$. It is worth noting that the corresponding circuit is significantly more convoluted, though, involving 42 gates for the random search compared to just 8 gates for the solution provided by the PPO agent. This non-optimal CHSH score is partially due to the fact that our agent is trained with modest computing power and does not rely on a large neural network. Therefore, the agent cannot learn complex circuit patterns required to reach higher CHSH scores. Moreover, learning a suboptimal policy has the additional negative effect of reducing the exploration in later episodes, which further reduces the chances to find highly non-trivial circuits leading to higher CHSH scores. A summary of the circuits found by the PPO agent and by random search for the different strategies is provided in Appendix˜B.
We now focus on the circuit depicted in Fig.˜1. This circuit appears as a good candidate for implementing a photonic Bell test with homodyne measurements. It has been found by the PPO agent interacting with a circuit of $N=4$ modes initialized with two two-mode squeezers and is composed of only two additional beam splitters. The third and fourth modes are used to herald the state using threshold detectors. This photonic circuit yields a CHSH score of $\mathcal{B}=2.068$ with optimal squeezing parameters corresponding to $3.9$ dB and $8\cdot 10^{-3}$ dB of squeezing, respectively. The heralding probability of the state is $3\cdot 10^{-6}$. Considering squeezers generating pulses at the MHz rate, the proposed circuit can prepare about $300$ states per second. Therefore, the statistics for a Bell test can be gathered in a realistic amount of time, e.g. $10^{6}$ shots in an hour.
Figure 2: CHSH score with respect to the distance between Alice, where state preparation occurs, and Bob. We consider an optical fiber with a loss of $0.2$ dB/km, and numerically optimized the circuits’ parameters for each step of $0.1$ km. As higher squeezing is beneficial for greater distance, we ensure squeezing of at most 1010dB to respect practical experimental range.
@@
To study the effect of photon loss, we assume that the optical circuit preparing and heralding the quantum state is located in Alice’s laboratory. Bob receives his part of the state, i.e the second mode, through an optical fiber link characterized by $0.2$ dB/km of attenuation. We assume that the optical signals are prepared at telecom wavelength, and, hence, that no conversions are necessary. In Fig.˜2, we plot the CHSH score as a function of the distance between Alice and Bob. For every value of distance, we re-optimize the parameters of the gates to maximize the CHSH score. In Table˜6 we provide the optimized value of these parameters with distance. We observe that a CHSH violation is still possible for distances up to $8.1$ km. As for the heralding probability of the state, this is not affected by photon loss occurring in Alice’s or Bob’s modes. See Section˜C.1 for more details on this analysis.
We then analyse the effect of imperfect detectors on circuit performance. Specifically, we assume that the threshold detectors used in the heralded modes have efficiency $\eta<1$. In Fig.˜3 we plot the heralding probability as a function of $\eta$, for fixed gate parameters optimal in the $\eta=1$ case (see Section˜C.2 for more details). We observe that the heralding probability decreases by at most one order of magnitude for efficiencies $\eta\geq 0.25$. Moreover, we observe that the efficiency of the detectors has barely any impact on the CHSH score. Specifically, from a value $\mathcal{B}=2.068$ at $\eta=1.0$, the CHSH score approaches $\mathcal{B}=2.067$ as $\eta\rightarrow 0$. This can be explained by the fact that the average number of photons reaching the threshold detectors is well below $1$. Indeed, if an inefficient threshold detector measures one mode out of a multi-photon entangled state, it detects each $n$-photon component of the state with a different probability, $1-(1-\eta)^{n}$, thereby modifying the photon number distribution of the unmeasured modes accordingly. This causes the conditional state of the unmeasured modes to depend, in principle, on the detector’s efficiency. However, if the multi-photon components in the measured mode are suppressed, the dependence of the conditional state on the detector efficiency becomes negligible, and the only real consequence is a reduction in the heralding probability.